OpenStudy (lexipoo1998):

Someone please help me! I have been asking these two questions all day and no ones seems to know how to do them. HELP!

OpenStudy (lexipoo1998):

Consider the function f(x) = the quantity x squared minus x minus 12 all over the quantity x plus 3. Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph.

OpenStudy (anonymous):

Ah, it's a simple graphing rational expressions problem.

OpenStudy (lexipoo1998):

I will give medal or whatever you want just please can you tell me what the answer is and how you got it?

OpenStudy (anonymous):

Begin by factoring the numerator.

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

You have x^2 -x -12 in the numerator. Its factors are (x-4)(x+3), but that is all divided by (x+3), so a factor cancels. When this occurs, you know your function has a hole when x = -3.

OpenStudy (lexipoo1998):

ok but how do I answer what it is asking?

OpenStudy (anonymous):

Plug that x value into the function, that is now f(x) = x-4, where x cannot be -3, to find the y value and the point of the discontinuity.

OpenStudy (anonymous):

This is part of the sol'n, lol, the point discontinuity or the "hole" in the function.

OpenStudy (lexipoo1998):

oh. so where are my intercepts?

OpenStudy (anonymous):

Your remaining function now is f(x) = x-4, which is a line. Your x-intercept can be found by making "y" equal to 0. Similarly, the y-intercept can be found by making the "x" value 0.

OpenStudy (anonymous):

This line, remember, has a hole at the point (-3, -7).

OpenStudy (lexipoo1998):

So thode would be the intercepts?

OpenStudy (lexipoo1998):

and wait would my point of discontinuity be -7?

OpenStudy (lexipoo1998):

@KinzaN are you still here?

OpenStudy (anonymous):

Yes, sorry. Well since you cancelled the factor (x+3) from both the numerator and the denominator, your hole or point discontinuity occurs when that factor is equal to 0, which is when x = -3. Plugging that value of x=-3 into your new function, y=x-4, results in an answer of -7. So your hole lies at the point f(-3)=-7 or (-3, -7).

OpenStudy (lexipoo1998):

wait so my point of discontinuity is -7 then what are my intercepts?

OpenStudy (lexipoo1998):

ugh this is confusing

OpenStudy (anonymous):

You find values on your x-axis--or x-intercepts--when your y-value is 0. So, 0= x-4 x=4 (4, 0). And, y= 0 - 4 y= -4 (0, -4).

OpenStudy (lexipoo1998):

ooooohhhhh ok. and I would describe the shape of the graph as a straight line right?

OpenStudy (anonymous):

A line with a slope of one and y-intercept of -4 with a hole at (-3, -7), I guess would be an awesome answer. :)

OpenStudy (lexipoo1998):

Lol ok. Can you help me with one more? I am about to post it.

OpenStudy (anonymous):

Haha, oookay. :) I'm avoiding my own homework by doing homework for others online, I've got time.

OpenStudy (lexipoo1998):

Lol ok THANK YOU!